# Ringdown amplitudes of nonspinning eccentric binaries

### v1

@@ Acknowledgements.
I am grateful to Mark Ho-Yeuk Cheung and Lionel London for constructive discussions, to Shilpa Kastha for stimulating interest in this problem and collaboration in the initial attempts to tackle it, to Jaime Redondo-Yuste, Vitor Cardoso, Rossella Gamba, Giada Caneva Santoro, Vasco Gennari, Michalis Agathos and Alex Nielsen for useful feedback on the manuscript, and to Manuela Campanelli, James Healy, Carlos Lousto and Yosef Zlochower for the creation and maintenance of the RIT public waveform database, on which this work is based. I thank the Institut des Hautes Études Scientifiques for its warm hospitality, where part of this work was developed. The present research was also partly supported by the “2021 Balzan Prize for Gravitation: Physical and Astrophysical Aspects”, awarded to Thibault Damour. I also thank Emanule Berti and the Bloomberg Center for Physics and Astronomy of The Johns Hopkins University, as well as Vitor Cardoso and the Centro de Astrofísica e Gravitação of Instituto Superior Técnico for generous hospitality during the last stages of this work. I acknowledge funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 847523 ‘INTERACTIONS’, and support from the Villum Investigator program by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101131233. Software: contents of this manuscript have been derived using the publicly available python software packages: cpnest, cython, matplotlib, numpy, pandas, pyRing, scipy, seaborn [79, 97, 98, 94, 99, 100, 31, 101, 102]. The obtained coefficients and a python implementation of the fits will be made publicly available at: github.com/GCArullo/noncircular_BBH_fits.

### v2

@@ Acknowledgements.
I am grateful to Mark Ho-Yeuk Cheung and Lionel London for constructive discussions, to Shilpa Kastha for stimulating interest in this problem and collaboration in the initial attempts to tackle it, to Jaime Redondo-Yuste, Vitor Cardoso, Rossella Gamba, Giada Caneva Santoro, Vasco Gennari, Michalis Agathos and Alex Nielsen for useful feedback on the manuscript, and to Manuela Campanelli, James Healy, Carlos Lousto and Yosef Zlochower for the creation and maintenance of the RIT public waveform database, on which this work is based. I thank the Institut des Hautes Études Scientifiques for its warm hospitality, where part of this work was developed. The present research was also partly supported by the “2021 Balzan Prize for Gravitation: Physical and Astrophysical Aspects”, awarded to Thibault Damour. I also thank Emanule Berti and the Bloomberg Center for Physics and Astronomy of The Johns Hopkins University, as well as Vitor Cardoso and the Centro de Astrofisica e Gravitação of Instituto Superior Técnico for generous hospitality during the last stages of this work. I acknowledge funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 847523 ‘INTERACTIONS’, and support from the Villum Investigator program by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101131233. Software: contents of this manuscript have been derived using the publicly available python software packages: cpnest, cython, matplotlib, numpy, pandas, pyRing, scipy, seaborn [81, 99, 100, 96, 101, 102, 31, 103, 104]. The obtained coefficients and a python implementation of the fits are publicly available at: github.com/GCArullo/noncircular_BBH_fits.
@@
Figure 9: Equal mass case for the dominant mode receiving mode-mixing, $(\ell,m)=(3,2)$. Top: fit in terms of a single variable representing noncircular initial conditions (“quasi-universal” case). Dots are colored by values of initial eccentricity. The spiral-like structure in the residuals indicates the degree of quasi-universality breaking for the quantity under consideration. Middle: fit in terms of the full two-degrees of freedom representing initial conditions. Dots are colored in terms of residuals. Bottom: histograms of residuals corresponding to the two-dimensional case.
@@
Figure 11: Parameter space (top), two-dimensional fits (middle) and three-dimensional fits (bottom) for the $(3,2,0)$ amplitude and phase parameters. The apparent spiky behaviour in the left central plot is a visualisation artifacts due to the surface curvature.